# MUSTAFAR: Promoting Unstructured Sparsity for KV Cache Pruning in LLM Inference

### v1

@@ Abstract
We demonstrate that unstructured sparsity significantly improves KV cache compression for LLMs, enabling sparsity levels up to 70% without compromising accuracy or requiring fine-tuning. We conduct a systematic exploration of pruning strategies and find per-token magnitude-based pruning as highly effective for both Key and Value caches under unstructured sparsity, surpassing prior structured pruning schemes. The Key cache benefits from prominent outlier elements, while the Value cache surprisingly benefits from a simple magnitude-based pruning despite its uniform distribution. KV cache size is the major bottleneck in decode performance due to high memory overhead for large context lengths. To address this, we use a bitmap-based sparse format and a custom attention kernel capable of compressing and directly computing over compressed caches pruned to arbitrary sparsity patterns, significantly accelerating memory-bound operations in decode computations and thereby compensating for the overhead of runtime pruning and compression. Our custom attention kernel coupled with the bitmap-based format delivers substantial compression of KV cache upto 45% of dense inference and thereby enables longer context length and increased tokens/sec throughput of upto 2.23$\times$ compared to dense inference. Our pruning mechanism and sparse attention kernel is available at https://github.com/dhjoo98/mustafar.
@@ 1 Introduction
In the age of Large Language Models (LLMs), advances in the machine learning domain [35, 2, 5] and the fast and efficient computing systems [18, 29] have lead to the emergence of highly capable LLMs that can summarize a book [19], write a compelling story [16], code a library [46], and generally reason over a longer context than ever before [6]. As LLMs are increasingly tasked with processing longer input sequences, the memory overhead associated with key-value (KV) caching has emerged as a critical bottleneck to scaling context length.
@@
Prior work has approached the challenge of KV cache memory overhead through techniques such as quantization [27, 14, 41, 45], low-rank approximation [40, 4, 31, 43, 23], token-wise eviction [44, 26, 22, 7, 1, 10], and structured pruning (e.g., channel-wise removal [38, 28]). The need to improve individual compression techniques has become increasingly important, especially as joint applications of multiple methods—such as pruning combined with token eviction [38], quantization with token-wise eviction [45], and low-rank approximation with quantization [4]—gain popularity. However, previous works on KV cache pruning have been limited to structured pruning, primarily due to the difficulty of efficiently leveraging finer-grained (i.e., unstructured) sparsity during execution. Effective pruning of the KV cache entails two core challenges: (1) achieving substantial reduction in KV cache size while preserving model accuracy, and (2) ensuring that the runtime pruning and compression processes are sufficiently efficient (i.e., the associated overhead must not outweigh the latency gains introduced by the resulting sparsity).
@@
Figure 1: High-level overview of Mustafar. Green region describes the pruning algorithm of Section 2, red region describes the custom sparse attention kernel of Section 3
@@
In this paper, we find that removing any constraint on the sparsity pattern, effectively unstructured sparsity can ensure that compressed KV cache perform with minimal model accuracy degradation while being pruned to a higher sparsity. In Section 2 (Green section of Figure 1), we first present our journey to find the optimal pruning algorithm for the key and value cache, based on the element magnitude distributions of the KV cache. We explore the feasibility of various pruning algorithms on both KV cache to conclude that applying a simple per-token magnitude-based pruning on both Key and Value cache is capable of preserving the model accuracy at a high sparsity, while also demonstrating strong compatibility with orthogonal compression techniques.
@@
Section 3 (Red section of Figure 1) discusses the next step: having induced sparsity in the KV cache, the challenge becomes leveraging the unstructured sparsity to reduce memory footprint and accelerate computation. To this end, we adopt a bitmap-based sparse format that serves both purposes. First, the bitmap enables maximal compression of matrices with arbitrary sparsity patterns. Second, this maximal compression of matrix operands translates into computational speedup of the attention operation, which is severely memory-bound on GPUs. Alongside the sparse format, we introduce the custom attention kernel tailored to operate on the bitmap-based sparse format. We see that the speedup of our attention kernel overshadows the latency introduced by runtime pruning and compression, meanwhile effectively compressing the KV cache to high sparsity with minimal accuracy degradation.
@@
In summary, we demonstrate that adopting unstructured sparsity in the KV cache, without imposing constraints on the pruning pattern, enables higher degrees of sparsity while preserving model accuracy. Furthermore, we introduce the necessary computational tools to support unstructured sparsity efficiently, ensuring that the derived high sparsity leads to gains in memory compression and increased end-to-end throughput in inference.
@@ 2 Pruning Algorithm for Unstructured Sparsity
We explore the potential unstructured sparsity on KV cache pruning by considering the two factors for Key and Value cache pruning: pruning direction and output-awareness. Pruning Direction refers to the axis along which sparsity is induced when selecting elements for removal. Since both the Key and Value caches are represented as matrices with dimensions $[tokens\times channels]$, we consider two primary pruning directions: per-channel pruning, which determines target sparsity across each channel (i.e., across tokens for each channel), and per-token pruning, which determines target sparsity across each token’s cache (i.e., across model dimensions for each token). Output-Awareness refers to the use of a scoring metric that serves as a proxy for estimating each element’s contribution to the operation’s output. Commonly employed in LLM weight pruning [32] and structured KV cache pruning [38], this technique involves computing a score for each pruning unit—such as a channel or an element—by taking the product of the corresponding element with its associated input. This approach effectively captures the element’s influence on the final output, guiding more informed pruning decisions. For a fair and effective comparison between pruning strategies, we uniformly employ a local dense window, where the recent 32 tokens remain untouched during the decode phase. Previous works [44, 38] have shown that this is effective in preserving model accuracy, meanwhile being small enough in size to not severely impact the compression.
@@ 2.1 Pruning Key Cache
In deciding the pruning direction, we build on top of the observation of KIVI [27], that Key cache exhibits distinct channel-wise outliers, where "channel" refers to the head dimension. (Figure 2(a)). This leads us to focus on per-token pruning for key cache, as it can effectively capture the elements in the outlier channel.
@@
Based on the same observation to perform structured pruning of individual channels, ThinK [38] incorporates output-awareness by using the a per-channel score of the accumulation of last 32 query, multiplied by each channel. To this end we compare the accuracy of ThinK [38], per-token magnitude-based unstructured pruning, and output-aware unstructured pruning of our design.
Figure 3: Per-token, output-aware pruning of Key cache.
Figure 3 elaborate the per-token output-aware unstructured pruning score of Key cache. The element-wise $L_{1}$ accumulation of the current and next 31 Query vector(blue) is multiplied element-wise across each token’s key vector(pink) to derive the pruning score(green). The absolute value of the score element in the corresponding position of each Key cache element is used to decide the elements to be pruned within a token’s Key vector. In other words, we formulate per-token output-aware unstructured pruning score $S$ of a Key cache $K$ to be:
@@
Table 1: Comparison of ThinK [38] structured pruning, per-token magnitude-based unstructured pruning, and per-token output-aware unstructured pruning on LongBench [3] with Llama-3-8B-Instruct Key cache. $K_{s}$ denotes Key cache sparsity
@@
ThinK (Structured) Unstructured Output-aware Unstructured Magnitude ThinK (Structured) Unstructured Output-aware Unstructured Magnitude
@@
In Table 1, we compare Llama-3-8B-Instruct accuracy of different pruning methods on LongBench [3]. For structured pruning, we see that even at a moderate sparsity, model accuracy retention is dismal compared to pruning to an unstructured sparsity pattern. Notably, unstructured pruning is capable of outperforming structured pruning even without the memory footprint of pruning scores involved with output-awareness. Applying output-awareness to unstructured pruning results in a slight improvement in the LongBench total average score, while performance across task averages is mixed — with each method outperforming the other on different tasks.
Key Cache Verdict: While “outlier channels” show promise for per-channel structured pruning, unstructured sparsity achieves higher accuracy at greater sparsity levels—even without output-awareness.
@@ 2.2 Pruning Value Cache
As shown in Figure 2(b), the Value cache exhibits a more uniform distribution of activations, making it challenging to apply the same channel-wise pruning without incurring a substantial degradation in model accuracy. This difficulty has led recent Value cache pruning approaches to be more susceptible to accuracy degradation.
Figure 4: Output-aware per-channel(red) and magnitude-based per-token pruning(pink) of Value cache. Magnitude-based per-token pruning is equivalent to output-aware per-token pruning(yellow).
With no discernible outliers in certain direction, we explore all possible combinations of (pruning direction, magnitude/output-aware) pairs. However, we are able to rule out per-token output-aware pruning, as the attention formulation $AttentionScore\times Value$ involves a multiply-and-accumulate operation along the token dimension. As seen in Figure 4, every element of a token’s Value cache is multiplied by the same element of the attention score, with each element’s impact on the output proportionate to the magnitude of each value. That is, for Value cache pruning, per-token magnitude-based pruning is already output-aware. For per-channel pruning, we prune each channel to the target sparsity in groups of 32 tokens, for compatibility with the local window size. For per-channel output-aware pruning, we accumulate the current and subsequent 31 attention score $\alpha$ of each token, which is then element-wise multiplied to the corresponding Value Cache ($V$) element. The following formula describe the pruning score $S$ of per-channel output-aware pruning:
$S=|V|\odot broadcast\left(\sum_{t=T}^{T+31}|\alpha_{t}|\right),\quad\text{where }\alpha_{t}\text{ is the attention score at time }t$
Table 2: Comparison of ThinK [38] structured pruning, per-channel magnitude-based unstructured pruning, per-channel output-aware unstructured pruning, and per-token magnitude-based pruning on LongBench [3] with Llama-3-8B-Instruct Value Cache. $V_{s}$ denotes Value cache sparsity
@@
ThinK (Structured) Magnitude (Per-channel) Output-aware (Per-channel) Magnitude (Per-token) ThinK (Structured) Magnitude (Per-channel) Output-aware (Per-channel) Magnitude (Per-token)
@@
As shown in the Table 2, we first see that applying structured pattern to Value cache pruning incurs significant accuracy degradation even in 50% sparsity. This is concurrent with ThinK [38] findings, which points to 30% sparsity as the upper-bound on ’acceptable’ accuracy. On the contrast, per-token magnitude pruning is capable of preserving model accuracy even at 70% sparsity. For per-channel pruning, we see that incorporating output-awareness boasts model accuracy retention almost to the level of per-token pruning. However we prefer per-token magnitude-based pruning for the following two reasons. First, output-aware per-channel value cache pruning requires access to the attention score which requires additional recomputation when used alongside FlashAttention [5], where the full attention score matrix does not materialize in the global memory. Second, per-token magnitude-based pruning allows smooth compatibility with orthogonal compression method token-wise eviction [21, 44], where the retained token’s KV cache can be pruned individually. We examine the accuracy of joint application in Section 4.2.
Value Cache Verdict: All unstructured pruning methods explored outperform structured pruning. Token-wise unstructured pruning, inherently output-aware by design, best preserves model accuracy even at high sparsity levels. While channel-wise pruning can achieve comparable accuracy with output-awareness, token-wise pruning offers advantages in both efficiency and modularity.
With the verdict in both Key and Value pruning, on Table 3 we finally validate the model accuracy retention of KV cache pruning in per-token magnitude-based fashion, when both Key and Value cache is pruned. Not only does unstructured sparsity enable value cache pruning to high sparsity, we see that both KV can be pruned to 70% sparsity while showing similar or better accuracy than Key-only 50% structured pruning of ThinK [38]. In Appendix A.1, methodology of this section is applied on Llama-2 7B in to reinforce the effectiveness of per-token magnitude-based pruning of KV cache.
@@ 3 Sparse Attention Kernel
Our findings establish that unstructured sparsity offers superior sparsity ratios over structured sparsity while preserving accuracy. A crucial contribution of Mustafar is to leverage this advantage to enable high compression efficiency, while minimizing the latency overhead of runtime pruning and compression. Prior compression methods—such as quantization, structured pruning, and token eviction—primarily reduce matrix dimensions or element bitwidths of the dense matrix products. In terms of efficiency, speedup from the reduced size of dense matrix products compensate for the additional latency introduced by compression (i.e. pruning score computation, quantization). In contrast, unstructured sparsity demands a fundamentally different approach.
Mustafar is motivated by the observation that attention operations in the autoregressive decode stage—namely, the Key $\times$ QueryT and Attention Score $\times$ Value computations—are matrix-vector products (MVs), significantly memory-bound on GPUs compared to the prefill stage.
(a) Mustafar attention kernel formulation.
(b) Bitmap-based sparse format.
@@
Figure 5: Overview of Mustafar attention fernel. In (a), multi-head, softmax, and normalization are omitted for simplicity.
To exploit this property, we introduce a bitmap-based sparse format as shown in Figure 5(b) to represent the pruned KV cache, thereby reducing data movement to GPU Streaming Multiprocessors. Pruning and compression are performed on-the-fly, with compression accelerated on GPU with a Triton kernel, and attention is computed directly on the compressed representation with a CUDA SpMV kernel.
@@
Figure 5(a) presents the Mustafar Sparse Attention Kernel. KV cache generated in prefill stage is pruned and compressed before the start of decode stage (therefore compatible with FlashAttention [5]). KV cache generated in decode stage is kept as-is (dense) while it is within the local window, then pruned and compressed afterwards. This entails the attention computations in the decode stage to be reformulated into two parts: SpMV for the compressed KV cache, and dense MV for the KV cache within the local window. Our bitmap-based format(Figure 5(b) consists of compressed tiles corresponding to a $1\times 64$ column of the pruned cache. Per-tile bitmap of 64 bits is used to represent the position of non-zeros, and tile offset is used to address the correct position of each tile’s starting non-zero.
@@
Mustafar SpMV kernel follows the load-as-compressed, compute-as-dense paradigm adopted by FlashLLM [37] and SpInfer [8], which target sparse matrix–dense matrix multiplication in LLM weight projection layers. The compressed KV cache is loaded from GPU global memory into registers in its compressed form, decompressed into shared memory, and then used for tile-wise dense computation. We evaluate the performance of the Mustafar attention kernel and quantify the runtime overhead of pruning and compression in Section 4.3. We further detail the formulation of the SpMV kernel, as well as the management of the compressed KV cache in Appendix C.

### v2

@@ Abstract
[…]
@@ 1 Introduction
In the age of Large Language Models (LLMs), advances in the machine learning domain [41, 2, 6] and the fast and efficient computing systems [21, 35] have led to the emergence of highly capable LLMs that can summarize a book [22], write a compelling story [18], code a library [53], and generally reason over longer contexts than ever before [7]. As LLMs are increasingly tasked with processing longer sequences, the memory overhead associated with key-value (KV) caching has emerged as a critical bottleneck to scaling context length.
@@
Prior work has approached the challenge of KV cache memory overhead through techniques such as quantization [30, 15, 48, 52], low-rank approximation [47, 4, 37, 50, 26], token-wise eviction [51, 29, 25, 8, 1, 11], and structured pruning (e.g., channel-wise removal [44, 31]). The need to improve individual compression techniques has become increasingly important, especially as joint applications of multiple methods, such as pruning combined with token eviction [44], quantization with token-wise eviction [52], and low-rank approximation with quantization [4], gain popularity. However, previous work on KV cache pruning have been limited to structured pruning, primarily due to the difficulty of efficiently leveraging finer-grained (i.e., unstructured) sparsity during execution. Effective pruning of the KV cache entails two core challenges: (1) achieving substantial reduction in KV cache size while preserving model accuracy, and (2) ensuring that the runtime pruning and compression processes are sufficiently efficient (i.e., the associated overhead must not outweigh the latency gains introduced by the resulting sparsity).
@@
Figure 1: High-level overview of Mustafar. Green region describes the pruning algorithm of Section 2, pink region describes the custom sparse attention kernel of Section 3.
@@
In this paper, we find that removing any constraint on the sparsity pattern, effectively unstructured sparsity can ensure that compressed KV cache perform with minimal model accuracy degradation while being pruned to a higher sparsity. In Section 2 (green region of Figure 1), we first present our journey to find the optimal pruning algorithm for the key and value cache, based on the element magnitude distributions of the KV cache. We explore the feasibility of various pruning algorithms on both KV cache to conclude that applying a simple per-token magnitude-based pruning on both Key and Value caches is capable of preserving the model accuracy at a high sparsity, while also demonstrating strong compatibility with orthogonal compression techniques.
@@
Section 3 (pink region of Figure 1) discusses the next step: having induced sparsity in the KV cache, the challenge becomes leveraging the unstructured sparsity to reduce memory footprint and accelerate computation. To this end, we adopt a bitmap-based sparse format that serves two purposes. First, the bitmap enables maximal compression of matrices with arbitrary sparsity patterns. Second, this maximal compression of matrix operands translates into computational speedup of the attention operation, which is severely memory-bound on GPUs. Alongside the sparse format, we introduce the custom attention kernel tailored to operate on the bitmap-based sparse format. We see that the speedup of our attention kernel overshadows the latency introduced by runtime pruning and compression, meanwhile effectively compressing the KV cache to high sparsity with minimal accuracy degradation.
@@
In summary, we demonstrate that adopting unstructured sparsity in the KV cache without imposing constraints on the pruning pattern enables higher degrees of sparsity while preserving model accuracy. Furthermore, we introduce the necessary computational tools to support unstructured sparsity efficiently, ensuring that the derived high sparsity leads to gains in memory compression and end-to-end inference throughput.
@@ 2 Pruning Algorithm for Unstructured Sparsity
We explore the potential unstructured sparsity on KV cache pruning by considering the two factors for Key and Value cache pruning: pruning direction and output-awareness. Pruning Direction refers to the axis along which sparsity is induced when selecting elements for removal. Since both the Key and Value caches are represented as matrices with dimensions $tokens\times channels]$, we consider two primary pruning directions: per-channel pruning, which determines target sparsity across each channel (i.e., across tokens for each channel), and per-token pruning, which determines target sparsity across each token’s cache (i.e., across model dimensions for each token). Output-Awareness refers to the use of a scoring metric that serves as a proxy for estimating each element’s contribution to the operation’s output. Commonly employed in LLM weight pruning [38] and structured KV cache pruning [44], this technique involves computing a score for each pruning unit such as a channel or an element by taking the product of the corresponding element with its associated input. This approach effectively captures the element’s influence on the final output, guiding more informed pruning decisions. For a fair and effective comparison between pruning strategies, we uniformly employ a local dense window, where the recent 32 tokens remain untouched during the decode phase. Previous works [51, 44] have shown that this is effective in preserving model accuracy, meanwhile being small enough in size to not severely impact the compression.
@@ 2.1 Pruning Key Cache
In deciding the pruning direction, we build on top of the observation of KIVI [30], that Key cache exhibits distinct channel-wise outliers, where "channel" refers to the head dimension (Figure 2(a)). This leads us to focus on per-token pruning for key cache, as it can effectively capture the elements in the outlier channel.
@@
Based on the same observation to perform structured pruning of individual channels, ThinK [44] incorporates output-awareness by using a per-channel score of the accumulation of last 32 query, multiplied by each channel. To this end we compare the accuracy of ThinK [44], per-token magnitude-based unstructured pruning, and output-aware unstructured pruning of our design.
Figure 3: Per-token, output-aware pruning of Key cache
Figure 3 elaborates the per-token output-aware unstructured pruning score of Key cache. The element-wise $L_{1}$ accumulation of the current and next 31 Query vector (blue) is multiplied element-wise across each token’s key vector (pink) to derive the pruning score (green). The absolute value of the score element in the corresponding position of each Key cache element is used to decide the elements to be pruned within a token’s Key vector. In other words, we formulate the per-token output-aware unstructured pruning score $S$ of a Key cache $K$ to be:
@@
Table 1: Comparison of ThinK [44] structured pruning, per-token magnitude-based unstructured pruning, and per-token output-aware unstructured pruning on LongBench [3] with Llama-3-8B-Instruct Key cache. $K_{s}$ denotes Key cache sparsity.
@@
ThinK
Unstructured
Output-aware
Unstructured
Magnitude
ThinK
Unstructured
Output-aware
Unstructured
Magnitude
@@
In Table 1, we compare Llama-3-8B-Instruct accuracy of different pruning methods on LongBench [3]. For structured pruning, we see that even at a moderate sparsity, model accuracy retention is dismal compared to pruning to an unstructured sparsity pattern. Notably, unstructured pruning is capable of outperforming structured pruning even without the memory footprint of pruning scores involved with output-awareness. Applying output-awareness to unstructured pruning results in a slight improvement in the LongBench total average score, while individual task performance is mixed with each method outperforming the other on different tasks.
Key Cache Verdict: While the existence of outlier channels with exceptionally high magnitudes show promise for per-channel structured pruning, unstructured sparsity achieves higher accuracy at greater sparsity levels, even without output-awareness.
@@ 2.2 Pruning Value Cache
As shown in Figure 2(b), Value cache exhibits more uniform distribution of activations, making it challenging to apply the same channel-wise pruning without incurring substantial degradation in model accuracy. This difficulty has led recent Value cache pruning approaches to be more susceptible to accuracy degradation.
Figure 4: Output-aware per-channel (red) and magnitude-based per-token (pink) pruning of Value cache. Magnitude-based per-token pruning is equal to output-aware per-token pruning (yellow).
With no discernible outliers in certain direction, we explore all possible combinations of (pruning direction, magnitude/output-aware) pairs. However, we are able to rule out per-token output-aware pruning, as the attention formulation $AttentionScore\times Value$ involves a multiply-and-accumulate operation along the token dimension. As seen in Figure 4, every element of a token’s Value cache is multiplied by the same element of the attention score, with each element’s impact on the output proportionate to the magnitude of each value. That is, for Value cache pruning, per-token magnitude-based pruning is already output-aware. For per-channel pruning, we prune each channel to the target sparsity in groups of 32 tokens, for compatibility with the local window size. For per-channel output-aware pruning, we accumulate the current and subsequent 31 attention score $\alpha$ of each token, which is then element-wise multiplied to the corresponding Value Cache ($V$) element. The following formula describes the pruning score $S$ of per-channel output-aware pruning:
[…]
Table 2: Comparison of ThinK [44] structured pruning, per-channel magnitude-based unstructured pruning, per-channel output-aware unstructured pruning, and per-token magnitude-based pruning on LongBench [3] with Llama-3-8B-Instruct Value Cache. $V_{s}$ denotes Value cache sparsity.
@@
ThinK
Magnitude
(Per-channel)
Output-aware
(Per-channel)
Magnitude
(Per-token)
ThinK
Magnitude
(Per-channel)
Output-aware
(Per-channel)
Magnitude
(Per-token)
@@
As shown in the Table 2, we first see that applying structured pattern to Value cache pruning incurs significant accuracy degradation even in 50% sparsity. This is concurrent with ThinK [44] findings, which points to 30% sparsity as the upper-bound on acceptable accuracy. In contrast, per-token magnitude pruning is capable of preserving model accuracy even at 70% sparsity. For per-channel pruning, we see that incorporating output-awareness boasts model accuracy retention almost to the level of per-token pruning. However, we prefer per-token magnitude-based pruning for the following two reasons. First, output-aware per-channel value cache pruning requires access to the attention score which requires additional recomputation when used alongside FlashAttention [6], where the full attention score matrix does not materialize in the global memory. Second, per-token magnitude-based pruning allows smooth compatibility with orthogonal compression method token-wise eviction [24, 51], where the retained token’s KV cache can be pruned individually. We examine the accuracy of joint application in Section 4.2.
Value Cache Verdict: All unstructured pruning methods explored outperform structured pruning. Among unstructured pruning methods, token-wise pruning, which is inherently output-aware by matrix multiplication formulation, best preserves model accuracy even at high sparsity levels. While channel-wise pruning with output-awareness can achieve comparable accuracy, token-wise pruning offers advantages in both efficiency and modularity.
With the two verdicts in Key and Value caches, on Table 3 we finally validate the model accuracy retention of per-token magnitude-based pruning with both Key and Value caches pruned. Not only can Value cache be pruned to high sparsity with unstructured sparsity, but both KV cache can be pruned to 70% sparsity while showing similar or better accuracy than Key-only 50% structured pruning of ThinK [44]. In Appendix A.1, methodology of this section is applied on Llama-2 7B to reinforce the effectiveness of per-token magnitude-based KV cache pruning.
@@ 3 Sparse Attention Kernel
Our findings establish that unstructured sparsity offers superior sparsity ratios over structured sparsity while preserving accuracy. In turn, a crucial contribution of Mustafar is to leverage this advantage to enable high compression efficiency while minimizing the latency overhead of runtime pruning and compression. Prior compression methods such as quantization, structured pruning, and token eviction reduce matrix dimensions or element bitwidths. In terms of efficiency, speedup from the reduced size of dense matrix operands compensates for the additional latency introduced by compression (i.e. pruning score computation, quantization). In contrast, unstructured sparsity with no regular reduction in dimensions or element bitwidth demands a different approach.
Mustafar is motivated by the observation that attention operations in the autoregressive decode stage, the Query $\times$ KeyT and Attention Score $\times$ Value computations are batch (different heads) of matrix-vector products (MVs) that are significantly memory-bound on GPUs compared to the prefill stage. To exploit this property, we extend the bitmap-based sparse format of Coruscant [20] as shown in Figure 5(a) to maximally compress the pruned KV cache. It consists of compressed tiles corresponding to a $1\times 64$ column of the pruned cache. Per-tile bitmap of 64 bits is used to represent the position of non-zeros, and tile offset is used to address the correct position of each tile’s starting non-zero. Pruning and compression are performed on-the-fly, with compression accelerated on GPU with a Triton kernel, and attention is computed directly on the compressed representation with a custom CUDA kernel that performs batch SpMV on the bitmap-based sparse format. Memory-bound decode-phase attention is accelerated by reducing the data movement from global memory to GPU Streaming Multiprocessors.
(a) Coruscant [20] bitmap-based sparse format
(b) Mustafar attention kernel formulation
@@
Figure 5: Overview of Mustafar sparse attention kernel. In (b), multi-head, softmax, and normalization are omitted for simplicity.
@@
Figure 5(b) and Algorithm 1 presents the Mustafar sparse attention kernel. KV cache generated in prefill stage is pruned and compressed before the start of decode stage, therefore compatible with prefill FlashAttention [6]. KV cache generated in decode stage is kept as-is (dense) while it is within the local window, then pruned and compressed afterwards. This entails the attention computations in the decode stage to be reformulated into two parts: SpMV for compressed KV cache (line 2 and 5 of Algorithm 1) and dense MV for the KV cache within the local window (line 1 and 5 of Algorithm 1).
Algorithm 1 Decode Phase Attention with Dense Local and Compressed KV Caches
1:Input: Query $\mathbf{Q}_{t}\in\mathbb{R}^{d}$; Local KV cache $\mathbf{K}_{L},\mathbf{V}_{L}\in\mathbb{R}^{d\times N_{d}}$, where $N_{d}$ is size of local window in tokens; Compressed KV cache $\mathbf{K}_{C},\mathbf{V}_{C}\in\mathbb{R}^{d\times N_{s}}$, where $N_{s}$ is number of compressed tokens.
2:Attention Score Computation
[…]
6:Output Computation
[…]
9:Return $\mathbf{O}_{t}$
@@
Mustafar SpMV kernel follows the load-as-compressed, compute-as-dense paradigm adopted by FlashLLM [43], SpInfer [9], and Coruscant [20], which target sparse matrix–dense matrix multiplication in LLM weight projection layers. The compressed KV cache is loaded from GPU global memory into registers in its compressed form, decompressed into shared memory, and then used for tile-wise dense computation. We evaluate the performance of the Mustafar attention kernel and quantify the runtime overhead of pruning and compression in Section 4.3. We further detail the formulation of the SpMV kernel, as well as the management of the compressed KV cache in Appendix C.